# Generalization of DNNs:

### v1

@@ 4.2 Breaking the Curse of Dimensionality with AccNets
Now that we know that composition of Sobolev functions can be easily learnable, even in settings where the curse of dimensionality should make it hard to learn them, we need to find a model that can achieve those rates. Though many models are possible  [^2] One could argue that it would be more natural to consider compositions of kernel method models, for example a composition of random feature models. But this would lead to a very similar model: this would be equivalent to a AccNet where only the $W_{\ell}$ weights are learned, while the $V_{\ell},b_{\ell}$ weights remain constant. Another family of models that should have similar properties is Deep Gaussian Processes [15]., we focus on DNNs, in particular AccNets. Assuming convergence to a global minimum of the loss of sufficiently wide AccNets with two types of regularization, one can guarantee close to optimal rates:
@@
Given a true function $f_{L^{*}:1}^{*}=f_{L^{*}}^{*}\circ\cdots\circ f_{1}^{*}$ going through the dimensions $d_{0}^{*},\dots,d_{L^{*}}^{*}$, along with a continuous input distribution $\pi_{0}$ supported in $B(0,b_{0})$, such that the distributions $\pi_{\ell}$ of $f_{\ell}^{*}(x)$ (for $x\sim\pi_{0}$) are continuous too and supported inside $B(0,b_{\ell})\subset\mathbb{R}^{d_{\ell}^{*}}$. Further assume that there are differentiabilities $\nu_{\ell}$ and radii $R_{\ell}$ such that $\left\|f_{\ell}^{*}\right\|_{W^{\nu_{\ell},2}(B(0,b_{\ell}))}\leq R_{\ell}$, and $\rho_{\ell}$ such that $Lip(f_{\ell}^{*})\leq\rho_{\ell}$. For an infinite width AccNet with $L\geq L^{*}$ and dimensions $d_{\ell}\geq d_{1}^{*},\dots,d_{L^{*}-1}^{*}$, we have for the ratios $\tilde{r}_{\ell}=\frac{\nu_{\ell}}{d_{\ell}^{*}+3}$:
@@
(1) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda R(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=\tilde{O}(N^{-\min\{\frac{1}{2},\tilde{r}_{1},\dots,\tilde{r}_{L^{*}}\}})$.
(2) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda\tilde{R}(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=\tilde{O}(N^{-\frac{1}{2}+\sum_{\ell=1}^{L^{*}}\max\{0,\tilde{r}_{\ell}-\frac{1}{2}\}})$.
There are a number of limitations to this result. First we assume that one is able to recover the global minimizer of the regularized loss, which should be hard in general [^3] Note that the unregularized loss can be optimized polynomially, e.g. in the NTK regime [28, 3, 16], but this is an easier task than findinig the global minimum of the regularized loss where one needs to both fit the data, and do it with an minimal regularization term. (we already know from [5] that this is NP-hard for shallow networks and a simple $F_{1}$-regularization). Note that it is sufficient to recover a network $f_{L:1}$ whose regularized loss is within a constant of the global minimum, which might be easier to guarantee, but should still be hard in general. The typical method of training with GD on the regularized loss is a greedy approach, which might fail in general but could recover almost optimal parameters under the right conditions (some results suggest that training relies on first order correlations to guide the network in the right direction [2, 1, 36]).
We propose two regularizations because they offer a tradeoff:
First regularization: The first regularization term leads to almost optimal rates, up to the change from $r_{\ell}=\frac{\nu_{\ell}}{d_{\ell}^{*}}$ to $r_{\ell}=\frac{\nu_{\ell}}{d_{\ell}^{*}+3}$ which is negligible for large dimensions $d_{\ell}$ and differentiabilities $\nu_{\ell}$. The first problem is that it requires an infinite width at the moment, because we were not able to prove that a function with bounded $F_{1}$-norm and Lipschitz constant can be approximated by a sufficiently wide shallow networks with the same (or close) $F_{1}$-norm and Lipschitz constant (we know from [5] that it is possible without preserving the Lipschitzness). We are quite hopeful that this condition might be removed in future work.
@@
The second and more significant problem is that the Lipschitz constants $Lip(f_{\ell})$ are difficult to optimize over. For finite width networks it is in theory possible to take the max over all linear regions, but the complexity might be unreasonable. It might be more reasonable to leverage an implicit bias instead, such as a large learning rate, because a large Lipschitz constant implies that the nework is sensible to small changes in its parameters, so GD with a large learning rate should only converge to minima with a small Lipschitz constant (such a bias is described in [26]). It might also be possible to replace the Lipschitz constant in our generalization bounds, possibly along the lines of [45].
Second regularization: The second regularization term has the advantage that it can be optimized with GD. Furthermore it does not require an infinite width, only a sufficiently large one. The issue is that it leads to rates that could be far from optimal depending on the ratios $\tilde{r}_{\ell}$: it recovers the same rate as the first regularization term if no more than one ratio $\tilde{r}_{\ell}$ is smaller than $\frac{1}{2}$, but if many of these ratios are above $\frac{1}{2}$, it can be arbitrarily smaller.
In Figure 2, we compare the empirical rates (by doing a linear fit on a log-log plot of test error as a function of $N$) and the predicted optimal rates $\min\{\frac{1}{2},\frac{\nu_{g}}{d_{in}},\frac{\nu_{h}}{d_{mid}}\}$ and observe a pretty good match.
Remark.
As can be seen in the proof of Theorem5, when the depth $L$ is strictly larger than the true depth $L^{*}$, one needs to add identity layers, leading to a so-called Bottleneck structure, which was proven emerge as a result of weight decay in [27, 26, 47]. These identity layers add a term that scales linearly in the additional depth $\frac{(L-L^{*})d_{min}^{*}}{\sqrt{N}}$ to the first regularization, and an exponential prefactor $(2d_{min}^{*})^{L-L^{*}}$ to the second.
@@ 5 Conclusion
We have given a generalization bound for Accordion Networks and as an extension Fully-Connected networks. It depends on $F_{1}$-norms and Lipschitz constants of its shallow subnetworks. This allows us to prove under certain assumptions that AccNets can learn general compositions of Sobolev functions efficiently, making them able to break the curse of dimensionality in certain settings, such as in the presence of unknown symmetries.
@@ Appendix A Experimental Setup
In this section, we review our numerical experiments and their setup both on synthetic and real-world datasets in order to address theoretical results more clearly and intuitively.
The code used for experiments is publicly available here.
@@ A.1 Dataset
A.1.1 Emperical Dataset
@@
We utilized four AMD Opteron 6136 processors (2.4 GHz, 32 cores) and 128 GB of RAM to generate our synthetic dataset. The maximum possible dataset size for 128 GB of RAM is approximately 52,500; however, we opted for a synthetic dataset size of 22,000 due to the computational expense associated with sampling the Matérn kernel. This decision was made considering the time complexity of $\mathcal{O}(n^{3})$and the space complexity of $\mathcal{O}(n^{2})$ involved. Out of the 22,000 dataset points, 20,000 were allocated for training data, and 2,000 were used for the test dataset
A.1.2 Real-world dataset: WESAD
In our study, we utilized the Wearable Stress and Affect Detection (WESAD) dataset to train our AccNets for binary classification. The WESAD dataset, which is publicly accessible, provides multimodal physiological and motion data collected from 15 subjects using devices worn on the wrist and chest. For the purpose of our experiment, we specifically employed the Empatica E4 wrist device to distinguish between non-stress (baseline) and stress conditions, simplifying the classification task to these two categories.
After preprocessing, the dataset comprised a total of 136,482 instances. We implemented a train-test split ratio of approximately 75:25, resulting in 100,000 instances for the training set and 36,482 instances for the test set. The overall hyperparameters and architecture of the AccNets model applied to the WESAD dataset were largely consistent with those used for our synthetic data. The primary differences were the use of 100 epochs for each iteration of Ni from Ns, and a learning rate set to 1e-5.
@@ A.2 Model setups
To investigate the scaling law of test error for our synthetic data, we trained models using $N_{i}$ datapoints from our training data, where $N=[100,200,500,1000,2000,5000,10000,20000]$. The models employed for this analysis included the kernel method, shallow networks, fully connected deep neural networks (FC DNN), and AccNets. For FC DNN and AccNets, we configured the network depth to 12 layers, with the layer widths set as $[d_{in},500,500,...,500,d_{out}]$ for DNNs, and $[d_{i}n,900,100,900,...,100,900,d_{out}]$ for AccNets.
To ensure a comparable number of neurons, the width for the shallow networks was set to 50,000, resulting in dimensions of $[d_{in},50000,d_{out}]$.
We utilized ReLU as the activation function and $L^{1}$-norm as the cost function, with the Adam optimizer. The total number of batch was set to 5, and the training process was conducted over 3600 epochs, divided into three phases. The detailed optimizer parameters are as follows:
@@
We conducted experiments utilizing 12 NVIDIA V100 GPUs (each with 32GB of memory) over a period of 6.3 days to train the synthetic dataset. In contrast, training the WESAD dataset required only one hour on a single V100 GPU.

### v2

@@ 4.2 Breaking the Curse of Dimensionality with AccNets
Now that we know that composition of Sobolev functions can be easily learnable, even in settings where the curse of dimensionality should make it hard to learn them, we need to find a model that can achieve those rates. Though many models are possible  [^2] One could argue that it would be more natural to consider compositions of kernel method models, for example a composition of random feature models. But this would lead to a very similar model: this would be equivalent to a AccNet where only the $W_{\ell}$ weights are learned, while the $V_{\ell},b_{\ell}$ weights remain constant. Another family of models that should have similar properties is Deep Gaussian Processes (Damianou & Lawrence, 2013)., we focus on DNNs, in particular AccNets. Assuming convergence to a global minimum of the loss of sufficiently wide AccNets with two types of regularization, one can guarantee close to optimal rates:
@@
Given a true function $f_{L^{*}:1}^{*}=f_{L^{*}}^{*}\circ\cdots\circ f_{1}^{*}$ going through the dimensions $d_{0}^{*},\dots,d_{L^{*}}^{*}$, along with a continuous input distribution $\pi_{0}$ supported in $B(0,b_{0})$, such that the distributions $\pi_{\ell}$ of $f_{\ell}^{*}(x)$ (for $x\sim\pi_{0}$) are continuous too and supported inside $B(0,b_{\ell})\subset\mathbb{R}^{d_{\ell}^{*}}$. Further assume that there are differentiabilities $\nu_{\ell}$ and radii $R_{\ell}$ such that $\left\|f_{\ell}^{*}\right\|_{W^{\nu_{\ell},2}(B(0,b_{\ell}))}\leq R_{\ell}$, and $\rho_{\ell}$ such that $Lip(f_{\ell}^{*})\leq\rho_{\ell}$. For an infinite width AccNet with $L\geq L^{*}$ and dimensions $d_{\ell}\geq d_{1}^{*},\dots,d_{L^{*}-1}^{*}$, we have for the ratios $\tilde{r}_{\ell}=\max\{\frac{d_{\ell-1}^{*}+3}{\nu_{\ell}},2\}$:
@@
(1) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda R(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=O(N^{-\frac{2}{2+\tilde{r}_{\max}}})$ for $\tilde{r}_{\max}=\max\{\tilde{r}_{1},\dots,\tilde{r}_{L}\}$.
(2) At a global min $\hat{f}_{L:1}$ of ~$\tilde{\mathcal{L}}_{N}(f_{L:1})+\lambda\tilde{R}(\theta)$, we have $\mathcal{L}(\hat{f}_{L:1})=O(N^{-\frac{2}{2+\tilde{r}_{sum}}})$ for $\tilde{r}_{sum}=2+\sum_{\ell}(\tilde{r}_{\ell}-2)$.
In the proof, we build a network $\hat{f}_{L:1}$ that approximates $f^{*}_{L^{*}:1}$ in the following manner: we adapt approximation results from (Bach, 2017) to approximate each Sobolev function $f^{*}_{\ell}$ by a shallow network $\hat{f}_{\ell^{\prime}}$ and if $L>L^{*}$ we add $L-L^{*}$ identity layers where the dimensionality $d^{*}_{\ell}$ is minimal (the parameter norm required to represent the identity on a $d$-dimensional representation is proportional to $d$, so it is optimal to add identity layers at the smallest dimension).
$F_{1}$-norm/Barron norm: This shows the power of replacing other parameter norm measures (such as the $(2,1)$-norm from Bartlett et al. (2017)) with the $F_{1}$-norm, by allowing us to repurpose the already existing approximation results for shallow networks. While we do not prove that a similar approximations cannot be achieved with small $(2,1)$-norm too, it seems unlikely, since generic Sobolev functions require a high variety of active neurons (a natural way to approximate them is to use a random feature approach (Rahimi & Recht, 2008)) which seems in contradiction with the $(2,1)$-norm which imposes a sparsity on the neurons.
Norm based vs Sparsity based: In (Schmidt-Hieber, 2020; Poggio & Fraser, 2024) a similar class of functions is considered and optimal rates $N^{-\frac{2}{2+r_{max}}}$ with $r_{max}=\max\{\frac{d_{0}^{*}}{\nu_{1}},\dots,\frac{d_{L^{*}-1}^{*}}{\nu_{L^{*}}}\}$ instead of $\tilde{r}_{max}=\max\{\frac{d_{0}^{*}+3}{\nu_{1}},\dots,\frac{d_{L^{*}-1}^{*}+3}{\nu_{L^{*}}},2\}$ are proven. The presence off the $+3$ comes from the shallow network approximation results and could potentially be removed. The bigger gap happens for easy tasks, where the ‘fast’ rates can be achieved $N^{-1}$ when $r_{max}\approx 0$, whereas our result can never predict rates faster than $N^{-\frac{1}{2}}$ because $\tilde{r}_{max}\geq 2$. This might be a fundamental gap, because norm-based complexity measures are known to be unable to yield fast rates (Srebro et al., 2010), so a control on the dimensionality of the model might be required to obtain these fast rates. Real world tasks are probably rarely easy enough to achieve rates faster than $N^{-\frac{1}{2}}$, e.g. LLMs scaling laws have been observed to be roughly $N^{-0.095}$ in (Kaplan et al., 2020).
Lipschitz constant vs operator norm: By evaluating the complexities $R,\tilde{R}$ on this approximation, we obtain two bounds illustrating the tradeoff between these two complexities. The first complexity measure based on Lipschitz constant $Lip(f_{\ell})$ leads to almost optimal rates, whereas the second complexity measure can lead to arbitrarily suboptimal rates when two or more ratios $\frac{d^{*}_{\ell-1}+3}{\nu_{\ell}}$ are larger than $2$. This is due to the fact to approximate a general Sobolev functions with a large ratio $\frac{d^{*}_{\ell-1}+3}{\nu_{\ell}}$ with a shallow network, the parameter norm $\|W_{\ell}\|_{F}^{2}+\|V_{\ell}\|_{F}^{2}$ needs to grow to infinity as $\epsilon\searrow 0$, leading to exploding operator norms, even though the Lipschitz constant remains bounded.
@@
Obviously, the Lipschitz constants $Lip(f_{\ell})$ are difficult to optimize over. For finite width networks it is in theory possible to take the max over all linear regions, but the complexity might be unreasonable. It might be more reasonable to leverage an implicit bias instead, such as a large learning rate, because a large Lipschitz constant implies that the network is sensible to small changes in its parameters, so GD with a large learning rate should only converge to minima with a small Lipschitz constant (such a bias is described in (Jacot, 2023b)). For this reason, we apply standard weight-decay to minimize the $F_{1}$-norms, and rely on the implicit bias to control the Lipschitz constants. Going further, techniques from (Wei & Ma, 2019) could be used to replace the Lipschitz constant in the bound with the maximum Jacobian over the dataset, which could then be optimized directly.
Impact of Depth: As can be seen in the proof of Theorem 5, when the depth $L$ is strictly larger than the true depth $L^{*}$, one needs to add identity layers, leading to a so-called Bottleneck structure, which was proven emerge as a result of weight decay in (Jacot, 2023a; b; Wen & Jacot, 2024). These identity layers add a term that scales linearly in the additional depth $\frac{(L-L^{*})d_{min}^{*}}{\sqrt{N}}$ to the first regularization, (and multiplicative factor of order $L$ to the second), see the proof for more details. The removal of the $2/3$ and $3/2$ exponents [^3] We achieve this by applying chaining accross the layers instead of applying to each layer separately. in our bounds in comparison to previous bounds Bartlett et al. (2017) allow us to get such a $O(L)$ bound instead of $O(L^{\frac{3}{2}})$. Interestingly, the first bound proposed in Golowich et al. (2020) to obtain a better depth dependence depends on the product of the Frobenius norms, which would grow exponentially as $\min_{\ell}\{d^{*}_{0},\dots,d^{*}_{L}\}^{(L-L^{*})}$ if the minimum dimension is 2 or more. Finally, note that by switching to a ResNet, there is no need for those identity layers, and the generalization becomes constant in depth.
Limitations: There are a number of limitations to this result. First we assume that one is able to recover the global minimizer of the regularized loss, which should be hard in general [^4] Note that the unregularized loss can be optimized polynomially, e.g. in the NTK regime (Jacot et al., 2018; Allen-Zhu et al., 2019; Du et al., 2019), but this is an easier task than finding the global minimum of the regularized loss where one needs to both fit the data, and do it with an minimal regularization term. (we already know from (Bach, 2017) that this is NP-hard for shallow networks and a simple $F_{1}$-regularization). Note that it is sufficient to recover a network $f_{L:1}$ whose regularized loss is within a constant of the global minimum, which might be easier to guarantee, but should still be hard in general. The typical method of training with GD on the regularized loss is a greedy approach, which might fail in general but could recover almost optimal parameters under the right conditions (some results suggest that training relies on first order correlations to guide the network in the right direction (Abbe et al., 2021; 2022; Petrini et al., 2023)).
Another limitation is that our proof requires an infinite width, because we were not able to prove that a function with bounded $F_{1}$-norm and Lipschitz constant can be approximated by a sufficiently wide shallow networks with the same (or close) $F_{1}$-norm and Lipschitz constant (we know from (Bach, 2017) that it is possible without preserving the Lipschitzness). We are quite hopeful that this condition might be removed in future work.
4.2.1 Experiments
We train our models on a synthetic dataset obtained by composing two Gaussian processes $Y=h(g(X))$ with Matérn kernels $K_{g},K_{h}$ chosen so that $g$ and $h$ have the right differentiability. In Figure 2, we compare the empirical rates (by doing a linear fit on a log-log plot of test error as a function of $N$) and rates $\min\{\frac{1}{2},\frac{2\nu_{g}}{2\nu_{g}+d_{in}},\frac{2\nu_{h}}{2\nu_{h}+d_{mid}}\}$ which seem to yield the best match. Note that this best fit for the rates is a mix of the theoretical lower bound $\min\{\frac{2\nu_{g}}{2\nu_{g}+d_{in}},\frac{2\nu_{h}}{2\nu_{h}+d_{mid}}\}$ and the upper bound $\min\{\frac{1}{2},\frac{2\nu_{g}}{2\nu_{g}+d_{in}+3},\frac{2\nu_{h}}{2\nu_{h}+d_{mid}+3}\}$, suggesting that the appearance of the $+3$ in the denominator might be an artifact of the proofs, whereas the $\frac{1}{2}$ might actually be optimal. But the amount of noise in our approximation of the rates does not allow us to give a definite answer for what the rates should be.
Notice also that all experiments were done with a simple $L_{2}$-regularization, which suggests the Lipshitz constants $Lip(f_{\ell})$ do not need to be optimized directly. The use of large learning rates and the edge of stability phenomenon (Cohen et al., 2021) could explain this implicit control of the Lipschitzness, because a large Lipschitz constant can lead to exploding gradients, and training with large learning rates could implicit avoid such parameters. Stability under large learning rate has been shown to control the regularity of the representations in previous works (Jacot, 2023a; b).
4.2.2 Computational graphs
Theorem 5 can be adapted to show that AccNets can also learn any ‘computational graph’ as in (Schmidt-Hieber, 2020; Poggio & Fraser, 2024), where a variables $v_{1},\dots,v_{K}$ each with their own dimension $d_{k}$. Each variable has a set of parents $P(k)\subset\{1,\dots,K\}$ on whom it depends $v_{k}=f_{k}((v_{m}:m\in P(k)))$ for a $\nu_{k}$-Sobolev or $F_{1}$ function $g_{k}$. We assume that the variables can be ordered in such a way that all parents of a variable belong have smaller indices, or equivalently that the directed graph with $K$ vertices, with edges connecting parents to their child is acyclic. The input variable is $1$ and output one is $K$.
It is then possible to organize the variables into ‘layers’ of variables that only have parents in the previous layers (possibly adding identities to keep certain variables from previous layers until they are needed). We say that the depth $L^{*}$ of the computational graph is the minimal number of layers required to find such a representation, and it matches the length of the longest directed path in the graph. Thus the complete function can be written as the composition of $L^{*}$ functions $f^{*}_{L:1}$, where each $f^{*}_{\ell}$ is the product of a few $g_{k}$s, e.g. $f^{*}_{\ell}(v_{1},v_{2},v_{3})=(g_{1}(v_{1},v_{2}),g_{2}(v_{3}),g_{3}(v_{1},v_{3}))$. The $F_{1}$-norm has the property that $F_{1}$-norm of the product of multiple functions is upper bounded by the sum of the $F_{1}$ norms, e.g. $\|f^{*}_{\ell}\|_{F_{1}}=\|g_{1}\|_{F_{1}}+\|g_{2}\|_{F_{1}}+\|g_{3}\|_{F_{1}}$, as well as the Lipshchitz constant $Lip(f^{*}_{\ell})=Lip(g_{1})+Lip(g_{2})+Lip(g_{3})$. One can then easily see that, assuming bounded Lipschitz constants, we can guarantee a rate of learning of $N^{-\frac{2}{2+\tilde{r}_{\max}}}$ for $\tilde{r}_{max}=\max_{k=1,\dots,K}\frac{3+\sum_{m\in P(k)}d_{k}}{\nu_{k}}$.
@@ 5 Conclusion
We have given a generalization bound for Accordion Networks and as an extension Fully-Connected networks. It depends on the $F_{1}$-norms and Lipschitz constants of its shallow subnetworks. This allows us to prove under certain assumptions that AccNets can learn general compositions of Sobolev functions efficiently, making them able to break the curse of dimensionality in certain settings, such as in the presence of unknown symmetries.
@@ Appendix A Experimental Setup
In this section, we review our numerical experiments and their setup both on synthetic and real-world datasets in order to address theoretical results more clearly and intuitively. The code used for experiments is publicly available here.
@@ A.1 Dataset
A.1.1 Empirical Dataset
By Corollary A.6 from (Tuo & Wu (2015)), the Reproducing kernel Hilbert space (RKHS) generated by Matérn kernel is norm equivalent to the Sobolev space $H^{\nu+d/2}(\Omega)$, so we utilized Matérn kernels to generate the synthetic dataset which retains Sobolev properties.
@@ A.1.1 Empirical Dataset
We used 128 GB of RAM to generate our synthetic dataset. Due to the Matérn Kernel’s time complexity of $\mathcal{O}(n^{3})$ and the space complexity of $\mathcal{O}(n^{2})$, the maximum possible dataset size for 128 GB of memory is approximately 52,500. Among the 52,500 dataset points, 50,000 were allocated for training, and 2,500 were used for the test dataset.
A.1.2 Real-world dataset: MNIST and WESAD
In our study, we utilized both MNIST (Modified National Institute of Standards and Technology) and WESAD (Wearable Stress and Affect Detection) to train our AccNets for classification tasks. As a standard benchmark for image classification tasks, MNIST consists of 60,000 grayscale images of 10 different handwritten digits. On the other hand, the WESAD dataset provides multimodal physiological and motion data collected from 15 subjects using devices worn on the wrist and chest. For the purpose of our experiment, we specifically employed the Empatica E4 wrist device to distinguish between non-stress (baseline) and stress conditions.
After preprocessing, the dataset comprised a total of 136,482 instances. We implemented a train-test split ratio of approximately 75:25, resulting in 100,000 dataset for the training set and the rest 36,482 dataset for the test set. The overall hyperparameters and architecture of the AccNets model applied to the WESAD dataset were largely consistent with those used for our synthetic data. The primary differences were the use of 100 epochs for each iteration of $N_{i}$ from $N$ in the model setup and a learning rate set to 1e-5.
@@ A.2 Model setups
For analyzing the scaling law of test error for our synthetic dataset, we trained models with gradually increasing $N_{i}$ data points from our training data, where $N=[100,200,500,1000,2000,5000,10000,20000,50000]$. The models utilized for our analysis are the kernel method, shallow networks, fully connected deep neural networks (FCNN), and AccNets. For FCNN and AccNets, we set the network depth to 12 layers, with the layer widths as $[d_{in},500,500,...,500,d_{out}]$ for DNNs, and $[d_{in},900,100,900,...,100,900,d_{out}]$ for AccNets.
In order to have a comparable number of neurons between the shallow networks and other deeper networks, the width for the shallow networks was set to 50,000, resulting in dimensions of $[d_{in},50000,d_{out}]$.
We utilized ReLU as the activation function and $L^{2}$-norm as the cost function, with the Adam optimizer. The total number of batch was set to 5, and the training process was conducted over 3600 epochs, divided into three phases. The detailed optimizer parameters are as follows:
@@
We conducted experiments utilizing 12 NVIDIA V100 GPUs (each with 32GB of memory) over approximately 7 days to train all the synthetic datasets. In contrast, training the WESAD and MNIST dataset required only one hour on a single V100 GPU.